OpenStudy (mendicant_bias):

Does anybody here know much about Perturbation Theory? I'm reading a book on it and am running into some issues. Also need some help with some plain old Calculus limits shown in the book that are confusing me.

OpenStudy (mendicant_bias):

\[z_{1}(\epsilon)=\lim_{\epsilon \rightarrow 0}\frac{1-\sqrt{1-\epsilon}}{\epsilon}=2z-1\]

OpenStudy (mendicant_bias):

\[z_{2}(\epsilon)=\lim_{\epsilon \rightarrow 0}\frac{1+\sqrt{1-\epsilon}}{\epsilon}=\infty\]

OpenStudy (mendicant_bias):

(I also don't honestly remember from limits how to mathematically verify that last one, either. Going to Wolfram|Alpha to poke around as well.)

OpenStudy (mendicant_bias):

Wolfram|Alpha gets something different, too; am I misreading the book? http://www.wolframalpha.com/input/?i=limit+as+epsilon+approaches+0+of+%281-sqrt%281-epsilon%29%29%2Fepsilon

ganeshie8 (ganeshie8):

*

OpenStudy (mendicant_bias):

What? But yeah, it's this part precisely right here that I don't get, math-wise. http://i.imgur.com/aSQN5rl.png

OpenStudy (mendicant_bias):

"The first quantitative step" is the beginning of the relevant section.

OpenStudy (mendicant_bias):

Hey, in that sentences where it says, "If epsilon equals zero, z_{1}(epsilon)=1/2", that's the exact same answer W|A got when taking the formal limit; so the other things aren't formal limits, or what?

OpenStudy (mendicant_bias):

I have no idea what the heck is going on in this book; this is literally the first few pages and there hasn't been any crazy math, and then this shows up and it just appears to not make any sense.

OpenStudy (mendicant_bias):

@tkhunny

OpenStudy (tkhunny):

#1 If it says \(\epsilon = 0\), it has nothing to do with a limit. Have you considered "rationalizing" your numerator?

OpenStudy (mendicant_bias):

Read the part beneath that where it declares epsilon zero, where it says "as epsilon approaches" and gives one of the results as infinity; that must be Calculus, or something else.

OpenStudy (mendicant_bias):

@tkhunny Right arrow, not equals sign. http://i.imgur.com/aSQN5rl.png

OpenStudy (tkhunny):

That makes a lot more sense. Now, how about the rationalizing?